Alexisbutton:

Courtney reads around 110 words per minute. She procrastinated and now has less than a week to finish her reading assignment. If the book she needs to finish has another 820 pages left and each page has around 350 words, how many hours should she set aside to complete her assignment show your work to receive full credit

jhonyy9:

given that Courtney reads around 110 words per minute using the rule of 3 simple you can calcule in how many minutes she reads one page so 350 words ,in this way you can writing 110 words reads in 1 min. 350 words reads in x min. cross multiplie and get x = 350/110 minutes = 35/11 min. so she read one page in 35/11 min. the book has 820 pages ,using the rule of 3 simple you can writing 1 page reads in 35/11 min. 820 pages reads in x min. x = 820*(35/11) min. what after make the calcule you need convert in hours